# When Experimental Economics Meets

### v1

@@ 3.1.3 Experimental Conditions
Building on the baseline condition, several variations are implemented to examine the validity of our tactics. In particular, we focus on the following three tactics which suggest that the corresponding variation may influence LLMs’ performance. Each condition includes 100 simulations for each model. SI Appendix, Section D provides detailed prompts for the baseline condition and these variations.
@@
    Persona (Tactic 2). We separately assign LLMs a range of demographics, including gender, age, education level, and ethnicity as in Chen et al. [23], as well as occupations as in Mei et al. [66]. [^4] We vary the persona by modifying the role specified in the system prompt, using statements such as “I want you to act as a male decision maker” and “I want you to act as a mathematician” [23]. For the variations of occupations, according to the method of Mei et al. [66], we additionally provide the description of a list of core and supplemental tasks associated with each occupation. Each specific persona corresponds to a unique condition.
@@
    Dialogue type (Tactic 6). Instead of using multi-turn dialogue, we consider a single-turn dialogue format that incorporates all 25 tasks in one prompt, instructing LLMs to make 25 decisions simultaneously.
@@
    Answer type (Tactic 7). We change the answer type from open-ended to multiple-choice [54, 21]. Specifically, the continuous budget set is discretized into 21 allocation options with a step size of 5 points, and LLMs are required to select one option for each task.
@@ 3.2.1 Rationality in Baseline Condition
Before discussing the economic rationality of LLMs, following Tactic 8, we first compute the proportion of invalid answers. Among the four models, GPT, DeepSeek and Qwen consistently produce valid answers across all 100 simulations in each domain. Only Llama has generated invalid answers in one simulation in the risk preference. These findings suggest that LLMs are generally capable of following instructions and providing the required outputs in budgetary decision tasks.
Next, we focus on the economic rationality of LLMs. The average CCEI scores of GPT, DeepSeek, Llama, and Qwen are 1.000, 1.000, 0.953, and 0.980 in the risk preference domain, and 0.994, 0.999, 0.968, and 0.994 in the social preference domain. [^5] To validate the sufficient power of detecting GARP violations, simulated CCEI is generated using random allocation across 25 decisions for 100 simulated agents [12]. The simulated CCEI is significantly lower than all four LLMs in both preference domains (all $p<0.01$). In human experiments using comparable budgetary tasks, the average CCEI ranges from 0.881 to 0.980 in the risk preference [24, 25, 47, 23, 21], and from 0.86 to 0.967 in the social preference [38, 39, 61, 60, 23]. Among the models evaluated, Llama and Qwen may fall within these ranges, while the other two models exhibit higher levels of rationality. We take human decisions in the baseline condition of Chen et al. [23] as a reference and depict the cumulative distribution of CCEI scores for human subjects and the four LLMs in Figure 1(a). The results indicate that GPT and DeepSeek outperform human subjects in both domains, Qwen surpasses humans in the social preference, while Llama performs worse than humans in the risk preference (all $p<0.01$, two-sided two-sample $t$-tests). [^6] We employ two-sided tests throughout all analyses. Unless otherwise specified, two-sample $t$-tests are used in this paper. Meanwhile, we conduct multiple hypotheses testing using the false discovery rate (FDR) method [5] in 8 tests (4 models $\times$ 2 preferences), and report corrected $p$ values. This approach is also applied in the comparison in decisions between LLMs and humans in Case Study 2 (Section 4.2.1), where we regard the 20 tests (4 models $\times$ 5 scenarios) as a family for the correction.
@@ 3.2.2 Rationality in Experimental Conditions
In the equation, $\mathcal{I}$ is the set of four LLMs, $\mathcal{M}$ denotes the set of preference domains in Case Study 1 or decision scenarios in Case Study 2. $\mathcal{S}$ represents the set of experimental conditions in each variation. $p$ denotes the $p$ value from the two-sample $t$-tests, and the significance level is set at 0.05.777Notably, we correct the $p$ values by controlling the false discovery rate (FDR) [5], where all tests within each variation are considered as a family for the correction. In Case Study 1, we perform multiple hypothesis tests in 112 tests (4 models $\times$ 2 preferences $\times$ 14 conditions) for the persona; while for the dialogue type and answer type, similar analysis is conducted in 8 tests (4 models $\times$ 2 preferences). In Case Study 2, the number of tests within a family is 280 (4 models $\times$ 5 scenarios $\times$ 14 conditions) for the persona and 20 (4 models $\times$ 5 scenarios) for the answer type. $\mathbb{1}(\cdot)$ is an indicator function. To further measure the sensitivity score under each preference domain (decision scenario) $m\in\mathcal{M}$, we also calculate $\lambda_{m}(\mathcal{S})=\Sigma_{i\in\mathcal{I},s\in\mathcal{S}}\frac{\mathbb{1}\{p_{ims}<0.05\}}{|\mathcal{I}||\mathcal{S}|}.$ By definition, both $\lambda$ and $\lambda_{m}$ belong to $[0,1]$. The higher the values, the larger the variation.
@@
Single-turn vs. multi-turn dialogue. Compared to the baseline condition, the single-turn dialogue approach does not substantially affect the CCEI scores of GPT (risk: 0.999 vs. 1.000, $p>0.1$, social: 1.000 vs. 0.994, $p<0.01$) and DeepSeek (risk: 1.000 vs. 1.000, social: 1.000 vs. 0.999, $p<0.01$) (Figure 2). However, when using the single-turn dialogue, the CCEI of Llama decreases from 0.953 to 0.841 in the risk preference and from 0.968 to 0.756 in the social preference; that of Qwen decreases from 0.980 to 0.739 in the risk preference and from 0.994 to 0.889 in the social preference (all $p<0.01$). As a result, $\lambda(\mathcal{S})$ equals 75.0%. These declines may arise from the limitation of single-turn dialogue in supporting sequential reasoning processes. Accordingly, the findings align with Tactic 6, supporting the adoption of a human-like multi-turn dialogue approach.
Open-ended vs. multiple-choice answer type. As shown in Figure 2, compared to the baseline condition using the open-ended answer type, the multiple-choice answer type does not affect the CCEI scores for GPT (risk: 1.000 vs. 1.000, social: 0.996 vs. 0.994, $p>0.1$) and DeepSeek (risk: 1.000 vs. 1.000, social: 0.994 vs. 0.999, $p=0.060$). In contrast, under the multiple-choice answer type, the average CCEI scores in both domains decrease significantly for Llama (risk: 0.853 vs. 0.953, $p<0.01$, social: 0.936 vs. 0.968, $p=0.033$) and Qwen (risk: 0.902 vs. 0.980, social: 0.916 vs. 0.994, both $p<0.01$). This decrease mirrors the pattern observed in GPT-3.5-Turbo [23]. Although the advancement from GPT-3.5-Turbo to GPT-4o over the past two years has eliminated the negative impact of discrete choices, this issue persists for smaller models such as Llama and Qwen (Tactic 7).
Other variations. In addition to the three variations discussed above, we also examine the influences of other tactics in SI Appendix, Section C. In summary, temperature settings between 0 and 1 (Tactic 1) do not affect the rationality of any of the four LLMs. The inclusion of examples (Tactic 5) does not affect the rationality in the risk preference, but does affect in the social preference. The existence of hypothetical incentives (Tactic 4) has no substantial impact on the rationality of four LLMs, while the rationality of Qwen surprisingly decreases as stake sizes increase in both preferences.
@@ 4.1 Experimental Design
We select the same four LLMs as in Case Study 1 and interact with them through the API in Python. The evaluated tasks include the dictator game, ultimatum game, public goods game, and bomb risk game. In each game, the LLMs are required to make one decision, with the following exception. In the ultimatum game, the LLMs separately play the role of proposer, who suggests an allocation, and responder, who determines the minimum proposal they are willing to accept. As a result, we collect the decisions of LLMs across five distinct decision scenarios from the four games. A detailed description of these tasks can be found in Mei et al. [66]. To avoid learning effects and simplify the analysis, we focus solely on the first round of decisions in each scenario. LLMs’ choices are required to be highlighted with “[[]]”. Data collection continues until 100 valid responses are obtained per scenario.
In the baseline condition, we adhere to the methodology from Case Study 1. Building upon this baseline, we implement two systematic variations concerning the persona and answer type. We do not examine the variation of dialogue type, as single-turn dialogue type is not feasible in these games. The prompts are detailed in SI Appendix, Section E.
@@
    Persona (Tactic 2). We independently vary gender, age, education level, and ethnicity following Chen et al. [23], as well as occupations based on Mei et al. [66].
@@
    Answer type (Tactic 7). Given that the feasible choice sets in all scenarios, except the public goods game, range from 0 to 100, we discretize this range into 21 options with a gap of 5. In the public goods game, where the feasible choice set ranges from 0 to 20, we maintain the same number of 21 options and reduce the gap to 1.
@@ 4.2.1 Preferences in Baseline Condition
    Public goods game. DeepSeek contributes more than human subjects in the public goods game, while Qwen contribute less (both $p<0.01$).
@@
Together, the four LLMs exhibit divergent preferences within each scenario, implying considerable heterogeneity in decisions between models. Moreover, Figure 3(a) shows that decisions from open-sourced models with smaller parameter sizes, such as Llama and Qwen, exhibit greater dispersion, which suggests higher heterogeneity in preferences within these models. In contrast, the behavior of GPT and DeepSeek is more consistent. [^8] For example, GPT consistently selects \$50 in the dictator game and sets the acceptance threshold at \$1 when acting as the responder in the ultimatum game. DeepSeek consistently contributes half of its endowment in the public goods game and opens half of the boxes in the bomb risk game. In a quantitative analysis, we calculate the normalized standard deviation of each LLM using $\tilde{\sigma}_{i}=\Sigma_{m\in\mathcal{M}}\frac{\sigma_{im}}{L_{m}|\mathcal{M}|}$. Here, $\sigma_{im}$ denotes the standard deviation of LLM $i$ in decision scenario $m\in\mathcal{M}$ in the baseline condition, and $L_{m}$ represents the length of the feasible set interval of scenario $m$, which is 20 for the public goods game and 100 for the other scenarios. The normalized standard deviations $\tilde{\sigma}_{i}$ for GPT, DeepSeek, Llama, and Qwen are 0.054, 0.030, 0.218, and 0.110, respectively. For comparison, the corresponding value for human decisions is 0.231.
@@ 4.2.2 Preferences in Experimental Conditions
Persona. Two key observations emerge. First, the decisions of LLMs in behavioral tasks are more sensitive to personas than in budgetary tasks ($\lambda(\mathcal{S})$: 35.7% vs. 0%, $p<0.01$, proportion test). Moreover, their choices in the public goods game are least likely to be influenced by personas among all games, with $\lambda_{m}(\mathcal{S})$ equal to 12.5%. Conversely, the sensitivity score significantly increases and reaches the highest when LLMs act as responders in the ultimatum game ($\lambda_{m}(\mathcal{S})$: 51.8% vs. 12.5%, $p<0.01$, proportion test). These findings suggest that the impact of personas depends on the nature of the tasks evaluated.
Second, LLMs’ outputs are more frequently influenced by occupational personas than by other demographic characteristics ($\lambda(\mathcal{S})$: 48.3% vs. 26.3%, $p<0.01$, proportion test). This underscores that the effect of personas is dependent on the specific types of personas. Taken together, these observations suggest that we cannot draw definitive conclusions about how personas influence experimental results, as their impact varies across tasks and persona types. These findings provide supporting evidence for Tactic 2, which argues against assigning specific personas in prompts, as the results may not be robust.
Open-ended vs. multiple-choice answer type. Compared to the open-ended answer type in the baseline condition, responding within a restricted set of options significantly alters LLMs’ outputs. Specifically, $\lambda(\mathcal{S})$ equals 55.0% under the multiple-choice answer type, indicating that the answer type significantly influences the decisions of more than half of the LLMs. Moreover, this value is marginally higher than that obtained under persona variations ($\lambda(\mathcal{S})$: 55.0% vs. 35.7%, $p=0.084$, proportion test), demonstrating that LLMs’ decisions are more sensitive to the answer type than personas.

### v2

@@ 3.1.3 Experimental Conditions
Based on the baseline condition, we implement several variations to investigate the impact of the considerations we have summarized. In particular, we focus on the following three factors which suggest that the corresponding variation may influence LLMs’ performance. Each condition includes 100 simulations for each model. Supplementary Section D provides detailed prompts for the baseline condition and these variations.
@@
    Persona. We separately assign LLMs a range of demographics, including gender, age, education level and ethnicity as in Chen et al. [23], as well as occupations as in Mei et al. [68]. [^4] We vary the persona by modifying the role specified in the system prompt, using statements such as “I want you to act as a male decision maker” and “I want you to act as a mathematician” [23]. For variations in occupations, according to the method of Mei et al. [68], we additionally provide the description of a list of core and supplementary tasks associated with each occupation. Each specific persona corresponds to a unique condition.
@@
    Dialogue type. Instead of using multi-turn dialogue, we consider a single-turn dialogue format that incorporates all 25 tasks in one prompt, instructing LLMs to make 25 decisions simultaneously.
@@
    Answer type. We change the answer type from open-ended to multiple-choice [55, 21]. Specifically, the continuous budget set is discretized into 21 allocation options with a step size of 5 points, and LLMs are required to select one option for each task.
@@ 3.2.1 Rationality in Baseline Condition
Before discussing the economic rationality of LLMs, we compute the proportion of invalid answers. Among the four models, GPT, DeepSeek and Qwen consistently produce valid answers across all 100 simulations in each domain. Only Llama generates invalid answers in one simulation in the risk preference. These findings suggest that LLMs are generally capable of following instructions and providing the required outputs in budgetary decision tasks.
Next, we focus on the economic rationality of LLMs. The average CCEI scores of GPT, DeepSeek, Llama, and Qwen are 1.000, 1.000, 0.953, and 0.980 in the risk preference domain, and 0.994, 0.999, 0.968, and 0.994 in the social preference domain. [^5] To validate the sufficient power of detecting GARP violations, simulated CCEI is generated using random allocation across 25 decisions for 100 simulated agents [12]. The simulated CCEI is significantly lower than all four LLMs in both preference domains (all $p<0.01$). In human experiments using comparable budgetary tasks, the average CCEI ranges from 0.881 to 0.980 in the risk preference [24, 25, 48, 23, 21], and from 0.860 to 0.967 in the social preference [39, 40, 63, 62, 23]. Among the models evaluated, Llama and Qwen may fall within these ranges, while the other two models exhibit higher levels of rationality. We take human decisions in the baseline condition of Chen et al. [23] as a reference and depict the cumulative distribution of CCEI scores for human subjects and the four LLMs in Figure 1(a). The results indicate that GPT and DeepSeek outperform human subjects in both domains, Qwen surpasses humans in the social preference, while Llama performs worse than humans in the risk preference (all $p<0.01$, two-sided two-sample $t$-tests). [^6] We employ two-sided tests throughout all analyses. Unless otherwise specified, two-sample $t$-tests are used in this paper. Meanwhile, we conduct multiple hypotheses testing using the false discovery rate (FDR) method [5] in 8 tests (4 models $\times$ 2 preferences), and report corrected $p$ values. This approach is also applied in the comparison in decisions between LLMs and humans in Case Study 2, where we regard the 20 tests (4 models $\times$ 5 scenarios) as a family for the correction.
@@ 3.2.2 Rationality in Experimental Conditions
In the equation, $\mathcal{I}$ is the set of four LLMs, $\mathcal{M}$ denotes the set of preference domains in Case Study 1 or decision scenarios in Case Study 2. $\mathcal{S}$ represents the set of experimental conditions in each variation. $p$ denotes the $p$ value from two-sample $t$-tests, and the significance level is set at 0.05.777Notably, we correct the $p$ values by controlling FDR [5], where all tests within each variation are considered as a family for correction. In Case Study 1, we perform multiple hypothesis testing in 112 tests (4 models $\times$ 2 preferences $\times$ 14 conditions) for the persona; while for the dialogue type and answer type, similar analysis is conducted in 8 tests (4 models $\times$ 2 preferences). In Case Study 2, the number of tests within a family is 280 (4 models $\times$ 5 scenarios $\times$ 14 conditions) for the persona and 20 (4 models $\times$ 5 scenarios) for the answer type. $\mathbb{1}(\cdot)$ is an indicator function. To further measure the sensitivity score under each preference domain (decision scenario) $m\in\mathcal{M}$, we also calculate $\lambda_{m}(\mathcal{S})=\Sigma_{i\in\mathcal{I},s\in\mathcal{S}}\frac{\mathbb{1}\{p_{ims}<0.05\}}{|\mathcal{I}||\mathcal{S}|}.$ By definition, both $\lambda$ and $\lambda_{m}$ belong to $[0,1]$. The higher the values, the higher the sensitivity.
@@
Single-turn vs. multi-turn dialogue. Compared to the baseline condition, the single-turn dialogue approach does not substantially affect the CCEI scores of GPT (risk: 0.999 vs. 1.000, $p>0.1$, social: 1.000 vs. 0.994, $p<0.01$) and DeepSeek (risk: 1.000 vs. 1.000, social: 1.000 vs. 0.999, $p<0.01$) (Figure 2). However, when using the single-turn dialogue, the CCEI of Llama decreases from 0.953 to 0.841 in the risk preference and from 0.968 to 0.756 in the social preference; that of Qwen decreases from 0.980 to 0.739 in the risk preference and from 0.994 to 0.889 in the social preference (all $p<0.01$). As a result, $\lambda(\mathcal{S})$ equals 75.0%. These declines may arise from the limitation of single-turn dialogue in supporting sequential reasoning processes.
Open-ended vs. multiple-choice answer type. As shown in Figure 2, compared to the baseline condition using the open-ended answer type, the multiple-choice answer type does not affect the CCEI scores for GPT (risk: 1.000 vs. 1.000, social: 0.996 vs. 0.994, $p>0.1$) and DeepSeek (risk: 1.000 vs. 1.000, social: 0.994 vs. 0.999, $p=0.060$). In contrast, under the multiple-choice answer type, the average CCEI scores in both domains decrease significantly for Llama (risk: 0.853 vs. 0.953, $p<0.01$, social: 0.936 vs. 0.968, $p=0.033$) and Qwen (risk: 0.902 vs. 0.980, social: 0.916 vs. 0.994, both $p<0.01$). This decrease mirrors the pattern observed in GPT-3.5-Turbo [23]. Although the advancement from GPT-3.5-Turbo to GPT-4o over the past two years has eliminated the negative impact of discrete choices, this issue persists for smaller models such as Llama and Qwen.
Other variations. In addition to the three variations discussed above, we also examine the influences of other factors in Supplementary Section C. In summary, temperature settings between 0 and 1 do not affect the rationality of any of the four LLMs. The inclusion of examples does not affect the rationality in the risk preference, but does affect in the social preference. The existence of hypothetical incentives has no substantial impact on the rationality of four LLMs, while the rationality of Qwen surprisingly decreases as stake sizes increase in both preferences.
@@ 4.1 Experimental Design
We select the same four LLMs as in Case Study 1 and interact with them through the API using Python. The evaluated tasks include the dictator game, the ultimatum game, the public goods game, and the bomb risk game. In each game, the LLMs are required to make one decision, with the following exception—in the ultimatum game, the LLMs separately play the role of proposer, who suggests an allocation, and responder, who determines the minimum proposal they are willing to accept. As a result, we collect the decisions of LLMs across five distinct decision scenarios from the four games. A detailed description of these tasks can be found in Mei et al. [68]. To avoid learning effects and simplify the analysis, we focus solely on the first round of decisions in each scenario. LLMs’ choices are required to be highlighted with “[[]]”. Data collection continues until 100 valid responses are obtained per scenario.
In the baseline condition, we adhere to the methodology in Case Study 1. Building upon this baseline, we implement two systematic variations concerning the persona and answer type. We do not examine the variation of dialogue type, as single-turn dialogue type is not feasible in these games. The prompts are detailed in Supplementary Section E.
@@
    Persona. We independently vary gender, age, education level, and ethnicity following Chen et al. [23], as well as occupations based on Mei et al. [68].
@@
    Answer type. Given that the feasible choice sets in all scenarios, except the public goods game, range from 0 to 100, we discretize this range into 21 options with a gap of 5. In the public goods game, where the feasible choice set ranges from 0 to 20, we maintain the same number of 21 options and reduce the gap to 1.
@@ 4.2.1 Preferences in Baseline Condition
    Public goods game. DeepSeek contributes more than human subjects in the public goods game, while Qwen contributes less (both $p<0.01$).
@@
Together, the four LLMs exhibit divergent preferences within each scenario, implying considerable heterogeneity in decisions between models. Moreover, Figure 3(a) shows that decisions from open-sourced models with smaller parameter sizes, such as Llama and Qwen, exhibit greater dispersion, which suggests higher heterogeneity in the inferred preferences of these models. In contrast, the behavior of GPT and DeepSeek is more consistent. [^8] For example, GPT consistently selects \$50 in the dictator game and sets the acceptance threshold at \$1 when acting as the responder in the ultimatum game. DeepSeek consistently contributes half of its endowment in the public goods game and opens half of the boxes in the bomb risk game. In a quantitative analysis, we calculate the normalized standard deviation of each LLM using $\tilde{\sigma}_{i}=\Sigma_{m\in\mathcal{M}}\frac{\sigma_{im}}{L_{m}|\mathcal{M}|}$. Here, $\sigma_{im}$ denotes the standard deviation of LLM $i$ in decision scenario $m\in\mathcal{M}$ in the baseline condition, and $L_{m}$ represents the length of the feasible set interval of scenario $m$, which is 20 for the public goods game and 100 for the other scenarios. The normalized standard deviations $\tilde{\sigma}_{i}$ for GPT, DeepSeek, Llama, and Qwen are 0.054, 0.030, 0.218, and 0.110, respectively. For comparison, the corresponding value for human decisions is 0.231.
@@ 4.2.2 Preferences in Experimental Conditions
Persona. Two key observations emerge. First, the decisions of LLMs in behavioral games are more sensitive to personas than in budgetary tasks ($\lambda(\mathcal{S})$: 35.7% vs. 0%, $p<0.01$, proportion test). Moreover, their choices in the public goods game are least likely to be influenced by personas among all games, with $\lambda_{m}(\mathcal{S})$ equal to 12.5%. Conversely, the sensitivity score increases significantly and reaches the highest when LLMs act as responders in the ultimatum game ($\lambda_{m}(\mathcal{S})$: 51.8% vs. 12.5%, $p<0.01$, proportion test). These findings suggest that the impact of personas depends on the nature of the tasks evaluated.
Second, LLMs’ outputs are more frequently influenced by occupational personas than by other demographic characteristics ($\lambda(\mathcal{S})$: 48.3% vs. 26.3%, $p<0.01$, proportion test). This underscores that the effect of personas is dependent on the specific types of personas. Taken together, these observations suggest that we cannot draw definitive conclusions about how personas influence experimental results, as their impact varies across tasks and persona types.
Open-ended vs. multiple-choice answer type. Compared to the open-ended answer type in the baseline condition, responding within a restricted set of options significantly alters LLMs’ outputs. Specifically, $\lambda(\mathcal{S})$ equals 55.0% under the multiple-choice answer type, indicating that the answer type significantly influences more than half of LLMs’ decisions. Moreover, this value is marginally higher than that obtained under persona variations ($\lambda(\mathcal{S})$: 55.0% vs. 35.7%, $p=0.084$, proportion test), demonstrating that LLMs’ decisions are more sensitive to the answer type than personas.
@@
5 Tactics for LLM Experiments
Drawing from key principles of human experiments, protocols in prior LLM research, and our own experimental findings, we present seven key tactics for conducting LLM experiments. These tactics serve two main purposes. First, they aim to align the procedures in LLM experiments with those in human experiments as closely as possible, thus facilitating the comparison in experimental results between LLM agents and human subjects. Second, with the burgeoning of LLM experiments, we hope these tips can be useful in improving the replicability of LLM experimental results. In the following, we will discuss each tactic by integrating insights from both experimental methods in economics and prompt engineering in artificial intelligence.
5.1 Experimental Design
Temperature. The selection of temperature involves the trade-off between adopting the solution with higher probability and exploring a broader solution space [81]. Human participants often exhibit a certain degree of heterogeneity. To mimic their behavior, we should allow the variation of responses between LLM agents. Meanwhile, both existing studies [23, 57, 76, 81] and our results (Supplementary Figure 3) have shown that there are no significant behavioral differences in LLMs’ outputs when different temperatures are used. Therefore, we suggest using the default temperatures recommended by model providers. When there are no default values, setting the temperature to 1 is a natural choice. This value does not scale the logits (see Supplementary Section B.3), thus best mirroring the application of softmax during the training process, which does not involve the temperature parameter [79]. In addition, some model providers allow users to specify a random seed parameter. To enhance the replicability of results, researchers can set the seed when prompting LLMs via the API. We summarize these discussions in the following.
Tactic 1: Use the default temperature and fix the random seed.
Persona. Following the conventions of experimental economics, where experimenters do not explicitly recruit participants with specific demographics unless this is part of the research design, we recommend not including specific personas in the prompts. Our two case studies further support this tactic that the use of personas may introduce uncontrolled variables without consistently yielding significant or generalizable effects. In general, LLMs should be regarded as counterparts to representative individuals from the general human population. However, if the goal is to assess whether LLMs with diverse demographic characteristics respond differently to tasks [23, 68, 90], we can implement it by assigning different personas. The caveat here is that the comparison in results between such studies should be cautious due to the differences in selected tasks, evaluated models, and persona types.
Incentive. Based on the induced value theory [85, 86], the use of monetary incentives has become a standard practice in laboratory experiments, although the extent to which human behavior differs under real versus hypothetical incentives remains inconclusive [e.g., 49, 58, 11, 44]. Since real monetary rewards cannot be actually applied to LLMs, the feasible alternative is to instruct LLMs to maximize their hypothetical payoffs. Therefore, clear payment rules could be included in the prompts as in the instructions for human participants. However, unlike humans, LLMs may lack an intrinsic utility function; their data generation is driven by externally imposed optimization rules in training processes. This indicates that LLMs are less likely to report untruthfully, the issue that incentives are designed to address. Note that although LLMs may exhibit hallucination characterized by nonsensical or unfaithful responses, this phenomenon results from the training data, training processes, and decoding strategies [52] rather than strategic deception. Based on these facts and the experimental results (Supplementary Figure 3), we conjecture that LLMs’ performance may change little when such hypothetical incentives are removed.
Tactic 3: Use the typical payment scheme to be comparable to human experiments.
5.2 Experimental Implementation
Tactic 4: Incorporate examples and comprehension questions especially for complex tasks.
Dialogue type. Since the structure of multi-turn dialogue is more similar to the sequential decision-making process of human subjects, we suggest using it for tasks involving multiple questions for a single agent. In contrast, although single-turn dialogue offers advantages in the efficiency of model inputs, it may constrain the economic rationality of certain models, particularly those with smaller parameter sizes (Figure 2).
Tactic 5: Use multi-turn dialogue for tasks with multiple questions.
Answer type. The construction of the answer format can influence model performance. For example, providing Likert scales or multiple choices may activate memory in the model training corpus, or enable models to infer experimental purposes and act in a demanding way [57]. Furthermore, discretizing continuous feasible sets has been shown to significantly decrease the performance of models in both Chen et al. [23] and our Case Study 1 (Figure 2). Consequently, the open-ended answer format is recommended rather than multiple-choice, unless the latter format is required by the nature of tasks. For example, the Linda problem requires multiple options with inclusion relationships to measure the conjunction fallacy [53].
Tactic 6: Use an open-ended answer format whenever feasible.
Invalid answer. In prior LLM experiments, the common practice is to exclude invalid answers from the dataset and continue data collection until the target sample size is reached [e.g., 23, 13, 45, 68]. However, the frequency of invalid answers also provides valuable information on model performance [66, 22]. Therefore, although the content of invalid answers may remain unused, we believe that the proportion of such responses should be documented and reported as a performance metric.
Tactic 7: Report the proportion of invalid responses.